OpenStudy (anonymous):

what are the factors of 1-27x^2

OpenStudy (anonymous):

\[(x - \frac{\sqrt{3}}{9})(x + \frac{\sqrt{3}}{9})\]

OpenStudy (anonymous):

\[1/3 (\sqrt3-9 x) (9 x+\sqrt3) \]

OpenStudy (anonymous):

\[(1-3x\sqrt{3})(1+3x\sqrt{3})\]

OpenStudy (anonymous):